# Source-Channel Separation Theorems for Distortion Perception Coding

### v1

@@ Appendix A Proof of Proposition 1
Consider two achievable distortion-perception pairs $(D_{1},P_{1})$ and $(D_{2},P_{2})$, for sufficiently large $n$ there exist codes $(f_{1},g_{1})$ and $(f_{2},g_{2})$ that achieve within $\epsilon$ of the distortion-perception pairs. We wish to show that the distortion perception pair $(\lambda D_{1}+(1-\lambda)D_{2},\lambda P_{1}+(1-\lambda)P_{2})$ is also achievable for any $\lambda\in(0,1)$. For any $\delta>0$, there exists a sufficient large integer $k_{1},k$, such that $|k_{1}/k-\lambda|\leq\delta$. We shall concatenate $k_{1}$ independent copies of $(f_{1},g_{1})$ and $k-k_{1}$ independent copies of $(f_{2},g_{2})$. This new code of length-$kn$ uses the channel $km$ times. The induced distortion is therefore upper-bounded by
@@
$\leq\lambda P_{1}+(1-\lambda)P_{2}+\epsilon+2\delta P_{\max},$ (35)
where in the first inequality we utilize the sub-decomposability of the perception measures, and $P_{\max}$ is the upper bound of the perception functions. By making $k$ and $n$ both large, $\epsilon$ and $\delta$ can both be made arbitrarily small. Therefore, there exists a code of length $n^{\prime}=kn$ that achieves within $\epsilon^{\prime}$ of the target $(D,P)$ pair.
We then show that as long as there exists a code function pair ($f,g)$ of length-$n$ that achieves within $\epsilon$ of the constraint $(D,P)$, codes exist for all sufficiently large $n^{\prime}$ to be within $2\epsilon$ of the constraint $(D,P)$. This can again be obtained by concatenating $k$ code copies of length-$n$ with a sufficiently large $k$, and some random segment of length-$n^{\prime\prime}$ to obtain a code of length-$(kn+n^{\prime\prime})$, where $0\leq n^{\prime\prime}<n$. When $k$ is sufficiently large, since both the distortion and perception measure functions are bounded and the perception measure function is sub-decomposable, the segment of length-$n^{\prime\prime}$ only induces negligible difference in the eventual distortion and perception. This completes the proof. ∎

### v2

@@ Appendix A Proof of Proposition 1
Consider two achievable distortion-perception pairs $(D_{1},P_{1})$ and $(D_{2},P_{2})$, for sufficiently large $n$ there exist codes $(f_{1},g_{1})$ and $(f_{2},g_{2})$ that achieve within $\epsilon$ of the distortion-perception pairs, each taking $m$ channel uses where $m\leq\kappa n$. We wish to show that the distortion perception pair $(\lambda D_{1}+(1-\lambda)D_{2},\lambda P_{1}+(1-\lambda)P_{2})$ is also achievable for any $\lambda\in(0,1)$. For any $\delta>0$, there exists a sufficient large integer $k_{1},k$, such that $|k_{1}/k-\lambda|\leq\delta$. We shall concatenate $k_{1}$ independent copies of $(f_{1},g_{1})$ and $k-k_{1}$ independent copies of $(f_{2},g_{2})$. This new code of length-$kn$ uses the channel $km$ times. The induced distortion is therefore upper-bounded by
@@
$\qquad\qquad\leq\lambda P_{1}+(1-\lambda)P_{2}+\epsilon+2\delta\max(P_{1},P_{2}),$ (35)
where we utilize the sub-decomposability of the perception measures in the first inequality. By making $k$ and $n$ both large, $\epsilon$ and $\delta$ can both be made arbitrarily small. Therefore, there exists a code of length $n^{\prime}=kn$ that achieves within $\epsilon^{\prime}$ of the target $(D,P)$ pair.
We then show that as long as there exists a code function pair ($f,g)$ of length-$n$ that achieves within $\epsilon$ of the constraint $(D,P)$, codes exist for all sufficiently large $n^{\prime}$ to be within $2\epsilon$ of the constraint $(D,P)$. This can again be obtained by concatenating $k$ code copies of length-$n$ with a sufficiently large $k$, and some random segment (with zero perception loss) of length-$n^{\prime\prime}$ to obtain a code of length-$(kn+n^{\prime\prime})$, where $0\leq n^{\prime\prime}<n$. When $k$ is sufficiently large, since the distortion function is bounded and the perception measure function is sub-decomposable, the segment of length-$n^{\prime\prime}$ only induces a negligible difference in the eventual distortion (and zero perception loss). This completes the proof. ∎